OpenStudy (anonymous):

Sqrt3(sqrt6+sqrt3)+(sqrt8-5) Solve in square root form please and thank you :3

OpenStudy (cristal):

Start by distributing. sqrt3*sqrt6 + sqrt3*sqrt3 +sqrt3= sqrt(3*6) + sqrt(3*3) + sqrt 3= sqrt(18) +3 + sqrt3. But we know sqrt18 can be sqrt(9*2)=sqrt9*sqrt2. So you're final answer should be: 3sqrt2+3+sqrt3